OpenStudy (aravindg):

If a,b,c,d are in G.P then (a+b+c+d)^2=?

OpenStudy (aravindg):

@ganeshie8

hartnn (hartnn):

does it lead to much simplification?? all you can use is ad = bc

hartnn (hartnn):

or maybe write, terms as a,ar, ar^2, ar^3 (b=ar, c=ar^2,d=ar^3)

OpenStudy (aravindg):

I have the answer it is (a+b)^2+(c+d)^2+2(b+c)^2

ganeshie8 (ganeshie8):

ahh that looks neat

hartnn (hartnn):

for the starters, just treat a+b as say A c+d as say B and expand (A+B)^2

hartnn (hartnn):

2AB= 2(a+b)(c+d) multiply them out and use the fact that ac = bd

hartnn (hartnn):

ad = bc***

OpenStudy (aravindg):

Tried it. But I am not able to progress.

hartnn (hartnn):

ok, ac+ad+bc+bd should equal (b+c)^2 ac +bc+bc+bd ----------------------------------------- so, ac+bd should equal b^2+c^2 we have a/b=b/c=c/d here ac= b^2 and bd = c^2 have a look :)

ganeshie8 (ganeshie8):

that gives below result : if a,b,c,d are in GP, then a+b, b+c, c+d also will be in GP

OpenStudy (aravindg):

Now that makes sense!! Thank you!! :)

hartnn (hartnn):

welcome ^_^